Knowing P = 56 kN and Q = 10 kN, determine the magnitude of the x-component of the resultant force or (Rx of the resultant) Note: if you solution indicates the resultant headed to the left, include the negative sign 45
Answer #1

56 kM = 39-72kN = (39.72-1,y) kN = 36,32 kry